Isayruiz:

OpenStudy (anonymous): triangle ABC is similar to triangle XYZ. If AB=4, and XY=10 and the area of triangle ABC is 7.8 units squared, find the area of triangle XYZ

Hero:

\(\triangle ABC \sim \triangle XYZ\) implies that \(\overline{AB}\) corresponds to \(\overline{XY}\), \(\overline{BC}\) corresponds to \(\overline{YZ}\), and \(\overline{AC}\) corresponds to \(\overline{XZ}\).

Isayruiz:

What's the answer can u help me

Hero:

Anyway, knowing that allows us to setup the following proportion: \(\dfrac{AB}{\text{area of }{ABC}} = \dfrac{XY}{\text{area of }{XYZ}}\)

Hero:

Plug-in the given information to the proportion and solve.

Isayruiz:

I don't know how to do it sadly I'm sorry

Hero:

when you plug in the given information you'll end up with: \(\dfrac{4}{7.8} = \dfrac{10}{x}\) where \(x\) represents the area of \(\triangle{XYZ}\)

Isayruiz:

19.5?

Hero:

Correct. The area of \(\triangle{XYZ}\) is 19.5 square units.

Isayruiz:

Thank u

Hero:

You're most welcome.

Hero:

@KittyNoir

Isayruiz:

One more

Hero:

You'll have to close this question first and re-post your question as a new question.